OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x − 10 y =2/3 x + 10 y = −2/3x − 10 y = −2/3x + 10

OpenStudy (anonymous):

the first one by the equation is y=2/3x-10

OpenStudy (anonymous):

What does that ".y = x-10" mean? Is It a typo?

OpenStudy (anonymous):

I think thats the first answer choice

OpenStudy (anonymous):

yeah its y=2/3x-10

OpenStudy (anonymous):

so@vinnv226 you think its the first one?

OpenStudy (anonymous):

hello?

OpenStudy (texaschic101):

2x -3y = 3....we want y = mx + b form and m will be the slope -3y = -2x + 3 y = -2/-3x + 3/-3 y = 2/3x - 1 the slope is 2/3 However you are looking for a perpendicular line....that means the slope is going to be the negative inverse of the slope....this means the slope you will need will be -3/2 now we have x = -8 and y = 2 and the slope(m) = -3/2 are you sure you wrote the problem correctly ? Because the slope for a perpendicular line is -3/2 and this does not match any of your answer choices. Now if it a parallel line, then that will work with the slope of 2/3.